OpenStudy (anonymous):

Richard’s average driving speed is 7 kilometers per hour faster than Gracie’s. In the same length of time it takes Richard to drive 292 kilometers, Gracie drives only 264 kilometers. What is Richard’s average speed?

OpenStudy (anonymous):

\[time=\frac{distance}{velocity} =\frac{292}{x+7}=\frac{264}{x}\] \[292x=264(x+7)..........292x-264x=1848..................x=1848/28=66 km/hr...........x+7=73km/hr\]

OpenStudy (anonymous):

these are my choices 73 km/hr 66 km/hr 59 km/hr 80 km/hr None of the above

OpenStudy (anonymous):

73 km/hr